### v1

# A model for Axial Non-Standard Interactions of neutrinos with quarks

Abstract
The neutrino oscillation experiments are setting increasingly strong upper bounds on the vectorial effective Non-Standard neutrino Interactions (NSI) with matter fields. However, the bounds on the axial NSI are more relaxed, raising the hope that studying the neutral current events at an experiment such as DUNE can give a glimpse of new physics. We build a model that gives rise to axial NSI with large couplings leading to observable deviation from the standard prediction at DUNE. The model is based on a $U(1)$ gauge symmetry with a gauge boson of mass $\sim 30$ GeV which can be discovered at the high luminosity LHC. Combining the LHC and DUNE discoveries, we can unravel the axial form of interaction. The cancellation of anomalies of the gauge group suggests new heavy quarks as well as a dark matter candidate. The new quarks mixed with the first generation quarks can also be discovered at the LHC. Moreover, they provide a seesaw mechanism that explains the smallness of the $u$ and $d$ quark masses. The dark matter has an axial coupling to the quarks which makes its discovery via spin dependent direct dark matter search experiments possible.
I Introduction
After establishment of the neutrino mass and mixing scheme as a solution for the solar and atmospheric neutrino anomalies, the main goal in this field has been determining the values of still unknown neutrino mass matrix parameters. In particular, to measure the Dirac CP violating phase of the PMNS neutrino mixing matrix, state-of-the-art long baseline experiments such as DUNE are under development. The precision of these experiments will enable us studying subdominant effects which might arise from new physics. Finding a footprint of new physics in the neutrino experiments is an opportunity that we do not want to miss. A wide class of new physics relevant for neutrino experiments can be described by the following effective potential
$\frac{G_{F}}{\sqrt{2}}[\bar{\nu}_{\alpha}\gamma^{\mu}(1-\gamma^{5})\nu_{\beta}][\bar{q}\gamma_{\mu}(\epsilon_{\alpha\beta}^{Vq}+\epsilon_{\alpha\beta}^{Aq}\gamma^{5})q]\ \ \ \ \ {\rm where}\ \ \ \ \ q\in\{u,d\}$ (1)
[…]
It will be dramatically interesting if new physics shows itself in the measurement of $\epsilon^{Au}$ and $\epsilon^{Ad}$ at DUNE. The aim of this paper is to build a model that leads to $\epsilon^{Vu}_{\alpha\beta},\epsilon^{Vd}_{\alpha\beta}\ll\epsilon^{Au}_{\tau\tau}\sim\epsilon^{Ad}_{\tau\tau}\sim O(1)$. We present a toy model based on a new gauge symmetry with a new gauge boson of mass $\sim 30$ GeV and axial coupling to the first generation quarks. For anomaly cancellation, we need to add new chiral fermions which will lead to interesting phenomenology, including a seesaw mechanism for the suppression of the $u$ and $d$ masses relative to the electroweak scale as well as a dark matter candidate with a mass of $\sim$ GeV which can be discovered by spin dependent direct dark matter search experiments.
This paper is organized as follows. In sect. II, with present the model and its field content and discuss the production and detection of its dark matter candidate. We also discuss bounds on the parameter space of the model. In sect. III, we discuss how various new particles introduced in this model can be discovered. In sect. IV, we first discuss generalization of the model to obtain an arbitrary lepton flavor structure for $\epsilon_{\alpha\beta}^{Au}=\epsilon_{\alpha\beta}^{Ad}$. We then suggest an alternative UV completion of the model without new colored particles. Our results are summarized in sect. V.
II The model
The effective interaction of form in Eq. (1) can be obtained by integrating out a massive gauge boson, $Z^{\prime}_{\mu}$, with axial couplings to the first generation quarks as well as couplings to neutrinos. As explained before, $\epsilon_{\tau\tau}^{Aq}$ is the least experimentally constrained element. In this section, we shall introduce a model for $|\epsilon_{\tau\tau}^{Aq}|\sim 1$. Let us introduce a gauge symmetry under which all three standard quark doublets have equal charges
$Q_{1}=\left(\begin{matrix}u_{L}\cr d_{L}\end{matrix}\right)\to\exp^{i\alpha}Q_{1}\ ,\ \ \ Q_{2}=\left(\begin{matrix}c_{L}\cr s_{L}\end{matrix}\right)\to\exp^{i\alpha}Q_{2}\ ,\ \ \ {\rm and}\ \ \ \ Q_{3}=\left(\begin{matrix}t_{L}\cr b_{L}\end{matrix}\right)\to\exp^{i\alpha}Q_{3}\ .$ (2)
Let us now assign opposite $U(1)$ charges to the first generation of the right-handed quarks
$u_{R}\to\exp^{-i\alpha}u_{R}\ \ \ {\rm and}\ \ \ \ d_{R}\to\exp^{-i\alpha}d_{R}\ .$
Then, the couplings of first generation quarks to $Z^{\prime}_{\mu}$ can be written as
$g_{Z}Z^{\prime}_{\mu}(\bar{u}\gamma^{\mu}\frac{1-\gamma_{5}}{2}u+\bar{d}\gamma^{\mu}\frac{1-\gamma_{5}}{2}d-\bar{u}\gamma^{\mu}\frac{1+\gamma_{5}}{2}u-\bar{d}\gamma^{\mu}\frac{1+\gamma_{5}}{2})=-g_{Z}Z^{\prime}_{\mu}(\bar{u}\gamma^{\mu}\gamma_{5}u+\bar{d}\gamma^{\mu}\gamma_{5}d)\ .$ (3)
We have therefore obtained axial form of coupling as we desired. Notice that our model respects the isospin symmetry. To cancel the $U(1)-SU(2)-SU(2)$ anomaly, we should assign $U(1)$ charges to lepton doublets, too. Since we want only $\nu_{\tau}$ to obtain effective coupling, we should assign a charge of $-9$ to $L_{3}$:
$L_{\tau}=\left(\begin{matrix}\nu_{\tau}\cr\tau_{L}\end{matrix}\right)\to\exp^{-9i\alpha}L_{\tau}\ \ \ \ {\rm under\ }U(1)\ .$ (4)
We then obtain
$\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}=\frac{9\sqrt{2}g_{Z}^{2}}{m_{Z^{\prime}}^{2}G_{F}}.$ (5)
As mentioned above, with this $U(1)$ charge assignments to the standard left-handed fermions, the $U(1)-SU(2)-SU(2)$ anomaly will cancel out. However, we should also ensure the cancellation of the $U(1)-SU(3)-SU(3)$ anomaly. Let us denote the $U(1)$ charges of $c_{R}$, $s_{R}$, $t_{R}$ and $b_{R}$ respectively with $\alpha_{c}$, $\alpha_{s}$, $\alpha_{t}$ and $\alpha_{b}$. It is straightforward to show that in order to ensure the $U(1)-SU(3)-SU(3)$ anomaly cancellation, $\alpha_{c}+\alpha_{s}+\alpha_{t}+\alpha_{b}=8$. The measured values of the Yukawa couplings of the SM Higgs to $\tau$, $b$ and $t$ are in remarkable agreement with the SM prediction Aad et al. (2024a). Thus, it seems the mass generation mechanism for the third generation fermions to be in fact the SM Higgs mechanism. Denoting the $U(1)$ charge of the Higgs with $\alpha_{H}$, this implies $\alpha_{\tau}={-9}-\alpha_{H}$, $\alpha_{b}=1-\alpha_{H}$ and $\alpha_{t}+\alpha_{b}=2$. Throughout this paper, we shall set $\alpha_{H}=0$ so we do not need to worry about the tree level decay modes such as $H\to Z^{\prime}Z^{\prime}$. The Yukawa couplings of the second generation quarks are less constrained by observation Aad et al. (2024a) so it is imaginable that they acquire mass through coupling to another Higgs doublet with a different $U(1)$ charge. For example, if we introduce a new Higgs that couples exclusively to $s_{R}$ and $Q_{2}$, we shall have the freedom to cancel the $U(1)-SU(3)-SU(3)$ anomaly. Another way to cancel the $U(1)-SU(3)-SU(3)$ anomaly is to add new colored particles charged under the new $U(1)$. The first case without new quarks is not necessarily more economic because mass generation for the first and second generation requires four new Higgs doublet with non-trivial $U(1)$ charges. Adding new colored particles makes the model more elegant and rich from the phenomenological point of view. In the following, we shall complete our model by adding new quarks. In sect. IV, we shall discuss the case without new quarks.
Let us introduce new quarks, $T_{L}$, $T_{R}$, $B_{L}$ and $B_{R}$ which are singlets of standard $SU(2)$ and are in the fundamental representation (triplet) of $SU(3)$ gauge group. The electric charges of $B$ (i.e., $B_{L}$ and $B_{R}$) and $T$ (i.e., $T_{L}$ and $T_{R}$) are respectively equal to $-1/3$ and $2/3$ as SM quarks. The field content of our model as well as their quantum numbers are shown in Table 1. The first and second generations of the leptons as well as the SM Higgs are taken to be neutral under the new $U(1)$ and are not shown in the table 1. Notice that $s_{R}$, $c_{R}$, $b_{R}$ and $t_{R}$ with charges opposite to those of $u_{R}$ and $d_{R}$ are taken to be mass eigenstates (rather than flavor eigenstates) to avoid Flavor Changing Neutral Currents (FCNC). We shall discuss more on FCNC in sect. III.
Table 1: Particles and their relevant quantum number under the gauge symmetry group. Here, T, D, and S stand for Triplet, Doublet, and Singlet respectively. The first and second generation leptons are not charged under new $\rm U(1)$ gauge and are not shown here.
Particles
$Q_{i}$ $t_{R},c_{R}$ $b_{R},s_{R}$ $u_{R}$ $d_{R}$ $L_{\tau}$ $\tau_{R}$ $T_{R}$ $T_{L}$ $B_{R}$ $B_{L}$ $\chi_{L}$ $\phi$ $\phi^{\prime}$
$\rm SU(3)$ T T T T T S S T T T T S S S
$\rm SU(2)$ D S S S S D S S S S S S S S
$\rm U_{Y}(1)$ $\frac{1}{6}$ $\frac{2}{3}$ $-\frac{1}{3}$ $\frac{2}{3}$ $-\frac{1}{3}$ $-\frac{1}{2}$1 $\frac{2}{3}$ $\frac{2}{3}$ $-\frac{1}{3}$ $-\frac{1}{3}$ 00 00 00
$\rm U(1)$ $1$ $1$ $1$ $-1$ $-1$ $-9$ $-9$ $1$ $-1$ $1$ $-1$ $9$ $-2$ $-18$
With the field content shown in table 1, all gauge anomalies cancel out. The left-handed $\chi_{L}$ is added to cancel the remaining $U(1)-U(1)-U(1)$ anomaly. Being neutral, $\chi_{L}$ can be a suitable DM candidate with a dazzling prospect of direct detection as we discuss below. Notice that with this charge assignment, $u$ and $d$ quarks cannot obtain mass via Yukawa coupling to the SM Higgs. In our model, their mass comes from mixing with the fourth generation quarks. Indeed, this model also explains the hierarchy between the first and third generation quarks as a bonus in an elegant fashion. Notice that we could put $(B_{L}\ T_{L})$ as well as $(B_{R}\ T_{R})$ in the $SU(2)$ electroweak doublets but the anomaly cancellation would then require smaller $U(1)$ charge for $L_{\tau}$, leading to smaller $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$.
The terms that lead to mass generation for the new fermions are
$\begin{split}&\lambda_{B}\phi\bar{B}_{L}B_{R}+\lambda_{T}\bar{T}_{L}T_{R}\phi+\frac{1}{2}\lambda_{\chi}\phi^{\prime}\chi_{L}^{T}c\chi_{L}+M_{Bd}\bar{B}_{L}d_{R}+M_{Tu}\bar{T}_{L}u_{R}+\lambda_{Bd}\bar{B}_{R}H^{\dagger}Q_{1}\\ &+\lambda_{Tu}\bar{T}_{R}H^{T}cQ_{1}+{\rm H.c.}\ \ {\rm where}\ \ c=\left(\begin{matrix}0&1\cr-1&0\end{matrix}\right).\end{split}$ (6)
We could add interactions to the second and third generation quarks as $\bar{B}_{R}H^{\dagger}Q_{2,3}$ and ${\bar{T}_{R}H^{T}c}Q_{2,3}$. For simplicity, we forbid these terms with an approximate global $U(1)$ symmetry under which only the second and third generation quarks are charged. Spontaneous symmetry breaking leads to masses for $\chi_{L}$ and the $Z^{\prime}$ boson as
$m_{\chi}=\lambda_{\chi}\langle\phi^{\prime}\rangle,\ \ {\rm and}\ \ m_{Z^{\prime}}=\sqrt{2}g_{Z^{\prime}}\left(4\langle\phi\rangle^{2}+324\langle\phi^{\prime}\rangle^{2}\right)^{1/2}$ (7)
and to the mass mixing of the first and fourth as
$(\bar{u}_{R}\ \bar{T}_{R})\left(\begin{matrix}0&M_{Tu}^{\star}\cr\lambda_{Tu}\langle H\rangle&\lambda_{T}\langle\phi\rangle\end{matrix}\right)\left(\begin{matrix}u_{L}\cr T_{L}\end{matrix}\right)+(\bar{d}_{R}\ \bar{B}_{R})\left(\begin{matrix}0&M_{Bd}^{\star}\cr\lambda_{Bd}\langle H\rangle&\lambda_{B}\langle\phi\rangle\end{matrix}\right)\left(\begin{matrix}d_{L}\cr B_{L}.\end{matrix}\right).$ (8)
Taking $M_{Tu}^{\star}$, $\lambda_{Tu}\langle H\rangle\ll\lambda_{T}\langle\phi\rangle$ and $M_{Bd}^{\star}$, $\lambda_{Bd}\langle H\rangle\ll\lambda_{B}\langle\phi\rangle$, the heavier mass eigenstates (i.e., the masses of the fourth generation quarks $T$ and $B$) will be equal to
$\ \ M_{B}=\lambda_{B}\langle\phi\rangle,\ \ {\rm and}\ \ M_{T}=\lambda_{T}\langle\phi\rangle\ .$
The lighter mass eigenvalues, which are the masses of the mass eigenstates $\tilde{u}$ and $\tilde{d}$, are respectively given by
$m_{u}=\frac{\lambda_{Tu}\langle H\rangle M_{Tu}^{\star}}{\lambda_{T}\langle\phi\rangle}=\frac{\lambda_{Tu}\langle H\rangle M_{Tu}^{\star}}{M_{T}}\ \ {\rm and}\ \ m_{d}=\frac{\lambda_{Bd}\langle H\rangle M_{Bd}^{\star}}{\lambda_{B}\langle\phi\rangle}=\frac{\lambda_{Bd}\langle H\rangle M_{Bd}^{\star}}{M_{B}}.$ (9)
With $M_{T}\gg M_{Tu},\lambda_{Tu}\langle H\rangle$ and $M_{B}\gg M_{Bd},\lambda_{Bd}\langle H\rangle$, there will be a seesaw mechanism that explains the smallness of the $u$ and $d$ masses relative to the electroweak scale. The mixing angles are as follows
${\rm Mixing\ between\ }u_{R}\ {\rm and}\ T_{R}=\frac{M_{Tu}^{\star}}{M_{T}}$ , ${\rm Mixing\ between\ }u_{L}\ {\rm and}\ T_{L}=\sin\alpha=\frac{\lambda_{Tu}\langle H\rangle}{M_{T}}$ (10)
${\rm Mixing\ between\ }d_{R}\ {\rm and}\ B_{R}=\frac{M_{Bd}^{\star}}{M_{B}}$ , ${\rm Mixing\ between\ }d_{L}{\rm and}\ B_{L}=\sin\beta=\frac{\lambda_{Bd}\langle H\rangle}{M_{B}}.$ (12)
Through the mixing, the mass eigenstate $\tilde{T}$ and $\tilde{B}$ can decay to the first generation quarks via decay modes such as $\tilde{T}\to W+\tilde{d}$ and $\tilde{B}\to W+\tilde{u}$. Since the $U(1)$ charges of $T_{R}$ and $B_{R}$ (of $T_{L}$ and $B_{L}$) are opposite to those of $u_{R}$ and $d_{R}$ (of $Q_{1}$), in the mass basis, $Z^{\prime}$ can have flavor changing couplings to the first and fourth generation quarks with a coupling given by $2g_{Z}$ times the relevant mixing. Both decay to $W$ and to $Z^{\prime}$ are suppressed by the mixing but since we expect $g_{Z}$ to be much smaller than the $SU(2)$ coupling, the decay to the $W$ boson will dominate. In our model, since the fourth generation left-handed quarks are $SU(2)$ singlets, their electroweak couplings are different from those of the first generation. As a result, the $Z$ boson can also have flavor changing coupling between left-handed first and fourth generation quarks leading to $T\to Zu$ and $B\to Zd$. The rates of these decay modes are comparable to the charged current ones. The bound on the masses of new quarks, which mainly decay into $W$ and first generation quarks, is Aad et al. (2024b)
$M_{T},M_{B}>1530~{}{\rm GeV}\ .$ (14)
Taking $\lambda_{T},\lambda_{B}$ in the perturbative range $(\lambda_{T},\lambda_{B}<4)$, Eqs. (5,7) yield
$\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}<0.9\ .$ (15)
As a result, our model can lead to $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}>0.1$ which can be probed by DUNE Abbaslu et al. (2024). Fig. 1 shows $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$ versus $m_{\chi}$ for different values of Yukawa coupling fixing $M_{T}$ and $M_{B}$ to 1550 GeV. For $m_{\chi}<10$ GeV, $\langle\phi\rangle\ll M_{T}/\lambda_{T},M_{B}/\lambda_{B}$ so the NSI coupling becomes independent of $m_{\chi}$. Notice that saturating the bound in Eq. (15) requires $9\langle\phi^{\prime}\rangle\ll\langle\phi\rangle$ (and therefore $m_{\chi}\ll M_{T},M_{B}$). As we shall see below, this is the range that we are interested in. Taking $M_{Z^{\prime}}>30$ GeV, we can safely use the four fermion effective interaction for neutrino scattering in experiments such as DUNE. Such large values of $m_{Z^{\prime}}$ requires $g_{Z}>0.02.$ On the other hand, to avoid the present bound Sirunyan et al. (2019), $g_{Z}$ should be smaller than 0.2.
Figure 1: $\epsilon_{\tau\tau}^{Au}$ and $\epsilon_{\tau\tau}^{Ad}$ versus $m_{\chi}$ for $\lambda_{\chi}=1$ and four different values: $\lambda_{B}=\lambda_{T}=4$ (dashed blue), $\lambda_{B}=\lambda_{T}=3$ (dashed purple), $\lambda_{B}=\lambda_{T}=2$ (dashed pink), and $\lambda_{B}=\lambda_{T}=1$ (dashed gray). We have fixed $M_{T}$ and $M_{B}$ to 1550 GeV.
Let us now discuss whether $\chi_{L}$ can play the role of Dark Matter (DM). To cancel the $U(1)-U(1)-U(1)$ anomaly, we had to add a SM singlet chiral field, $\chi_{L}$. In principle, our gauge group allows a coupling of form $\bar{\chi_{L}^{c}}H^{T}cL_{\tau}$ which leads to the decay of $\chi_{L}$ to $\nu_{\tau}$ plus muon or $s$-quarks pairs. Imposing a $Z_{2}$ symmetry under which only $\chi_{L}$ is odd this term will be forbidden and $\chi_{L}$ can be stable. Integrating out the intermediate $Z^{\prime}$, the effective coupling between $\chi$ and the first generation quarks can be written as
$\frac{9g_{Z}^{2}}{m_{Z^{\prime}}^{2}}\left(\bar{\chi}\gamma^{\mu}(\frac{1-\gamma_{5}}{2})\chi\right)\left(\bar{q}\gamma_{\mu}\gamma_{5}q\right).$ (16)
This effective interaction leads to spin-dependent interaction between dark matter and the nucleons (proton and neutron) with the following cross section Cirelli et al. (2024):  .
$\sigma_{SD}^{N}\simeq 20.97\left(\frac{m_{N}m_{\chi}}{m_{N}+m_{\chi}}\right)^{2}\frac{g_{Z}^{4}}{m_{Z^{\prime}}^{4}}=5.24\left(\frac{m_{N}m_{\chi}}{m_{N}+m_{\chi}}\right)^{2}\frac{1}{\Big{(}(2M_{T}/\lambda_{T})^{2}+(18m_{\chi}/\lambda_{\chi})^{2}\Big{)}^{2}}\ .$ (17)
Notice that our model is isospin invariant, predicting equal cross sections for scattering off the neutron and the proton. Fig 2 shows the prediction of our model for the scattering cross section for different values of $\lambda_{T}=\lambda_{B}$ along with the bounds from the spin dependent direct detection. As seen from these figures, the direct dark matter detection bounds from the PICO experiment Amole et al. (2019) forces $\chi$ to be lighter than a few GeV:
$m_{\chi}<2~{}{\rm GeV}\ \ {\rm for}\ \lambda_{T}=\lambda_{B}=4\ .$ (18)
Taking $m_{Z^{\prime}}>30$ GeV, we can therefore use the four Fermi effective formalism to compute the cross section of the annihilation of dark matter pair to the standard model fermions Beltran et al. (2009). The cross section of the annihilation to the SM fermions for $\langle\phi\rangle>400$ GeV (i.e., $M_{T},M_{B}>1600$ GeV) will be much smaller than 1 pb, satisfying the bounds from the indirect dark matter searches Beltran et al. (2009). However, if the reheating temperature is higher than the mass of $\chi$ and $Z^{\prime}$, they will come into thermal equilibrium with the plasma in the early universe. An over-abundance of DM can be avoided if the reheating temperature is low Boddy et al. (2024). To have freeze-out mechanism with the annihilation cross section $\sigma_{ann}\sim 1$ pb, we need to turn on new interaction(s). The $\phi^{\prime}$ particle can be a $s$-channel mediator for this annihilation. For example, if $\phi^{\prime}$ mixes with the neutral component of a $SU(2)$ triplet with the same $U(1)$ charge, we can obtain annihilation into neutrinos, $\chi\chi\to\nu_{\tau}\nu_{\tau}$, which is less constrained by indirect dark matter searches than annihilation to quarks or tau. Another option is a coupling of $\phi^{\prime}$ to two light singlet scalars, $a$. This can take place via a quartic coupling of form $\lambda_{a\phi^{\prime}}|\phi^{\prime}|^{2}a^{2}$. The annihilation cross section will be given by
$\sigma v=\frac{\lambda_{a\phi^{\prime}}^{2}m_{\chi}^{2}}{128\pi(m_{\phi^{\prime}}^{2}-4m_{\chi}^{2})^{2}}\sqrt{1-\frac{m_{a}^{2}}{m_{\chi}^{2}}}\Big{(}1+\mathcal{O}(v^{2})\Big{)},$ (19)
which can be equal to 1 pb for $m_{\phi^{\prime}}\sqrt{1-4m_{\chi}^{2}/m_{\phi^{\prime}}^{2}}=31~{}{\rm GeV}\sqrt{\lambda_{a\phi^{\prime}}m_{\chi}/{\rm GeV}}$. If $m_{\chi}\sim$GeV, the freeze-out will take place around the time of QCD phase transition so the density of $a$ at the neutrino decoupling era will be suppressed and it will contribute less than 1 to the relativistic degrees of freedom ($\Delta N_{eff}$) at the neutrino decoupling era and later. At temperatures below $m_{\chi}/20$, the $a$ particles will decouple from the plasma and their density will be entropy suppressed. Taking the $a$ mass smaller than $\sim 100$ eV, their contribution to the matter budget of the universe today will be negligible, making $\chi_{L}$ the prime dark matter candidate.
The $\chi$ particle can scatter off each other via $Z^{\prime}$ and/or $\phi^{\prime}$ exchange but the cross section will be much below values that can have observable effects Beltran et al. (2009): $\sigma(\chi\chi\to\chi\chi)/m_{\chi}\ll 0.1~{}{\rm cm}^{2}/{\rm gr}.$ As a result, our dark matter candidate can be regarded as collision-less dark matter for structure formation and galaxy cluster collisions.
Figure 2: Spin-dependent DM-nucleon scattering cross section vs. DM mass for (a) protons and (b) neutrons. The solid curves are the present bounds from spin dependent direct dark matter search experiments and the dashed curves are the predictions of our model. The magenta solid line represents PICO-60 Amole et al. (2019), the red solid line shows LZ 2022 Aalbers et al. (2023), the green solid line depicts XENONnT 2023 Aprile et al. (2023), the brown solid line demonstrates PandaX-II Xia et al. (2019), the orange dashed line represents LUX-2017 Aalbers et al. (2023), and the cyan solid line shows LUX-2013 Akerib et al. (2016). The spin dependent cross section $\sigma^{N}_{SD}$ is plotted for $\lambda_{\chi}=1$ and four different values: $\lambda_{B}=\lambda_{T}=4$ (dashed blue), $\lambda_{B}=\lambda_{T}=3$ (dashed purple), $\lambda_{B}=\lambda_{T}=2$ (dashed pink), and $\lambda_{B}=\lambda_{T}=1$ (dashed gray), fixing $M_{T}=M_{B}=1550$ GeV.
III Testable signals of the model
In the previous section, we found that in our model the axial NSI coupling, $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$, can be as large as $U(1)$ which as shown in Abbaslu et al. (2024) will have observable impact on the number of neutral current events at the far detector of DUNE. This coupling can also affect the diffusion of the tau neutrinos out of the supernova core Cerdeño et al. (2023); Fiorillo et al. (2024); Cerdeño et al. (2021). In fact, the main goal of the present model is to obtain axial NSI with large couplings. To build a viable anomaly free gauge interaction that leads to such NSI, we introduce new particles, $Z^{\prime}$, heavy quarks ($T$ and $B$) and the dark matter candidate ($\chi_{L}$). An exciting aspect of our model is that each of these particles may show up in the upcoming experiments, making our toy model “super-testable.” Let us discuss their potential signatures one by one.
Discovery of $Z^{\prime}$: A low mass gauge boson such as $Z^{\prime}$ coupled to $u$ and $d$ can be discovered as a resonance in the dijet distribution in the proton proton colliders. In fact, the LHC data with an integrated luminosity of 35.9 fb-1 has already set a bound of $\sim 0.2$ on the coupling of $Z^{\prime}$ with a mass of $\sim 30$ GeV to the quarks Sirunyan et al. (2019). The high luminosity phase of the LHC with 3 ab-1 integrated luminosity (100 times more data) is expected to probe couplings one order of magnitude lower, delving into the most interesting range of the parameter space of this model. However, this discovery cannot determine whether the $Z^{\prime}$ coupling to the quarks is vectorial or axial. Observing a deviation of the number of the neutral current events at the far detector of DUNE, along with the stringent bounds on $\epsilon^{Vq}$ from the other neutrino experiments, can testify the axial form of the interaction.
Signatures of $\chi_{L}$: As we saw in the previous section, this model naturally offers a candidate for a dark matter which can undergo spin dependent dark matter scattering off nuclei. Direct dark matter search experiments already constrain dark matter mass to be lighter than $O(1)$ GeV within this model with $\epsilon^{Au}=\epsilon^{Ad}\sim 1$. The upcoming direct dark matter searches can find $\chi_{L}$. The indirect dark matter search bounds from dwarf galaxies are satisfied in our model. If the production mechanism for $\chi_{L}$ in the early universe is freeze-out with annihilation to dark sector, we would expect fractional contribution to extra relativistic degrees of freedom in the early universe which can be eventually tested by improvements on the BBN and CMB measurements.
Discovery of $T$ and $B$: As emphasized in the previous section $t_{R}$, $c_{R}$, $b_{R}$ and $s_{R}$ with $U(1)$ charge of 1 are taken to be mass eigenstates. However, mass eigenstates $\tilde{u}_{R}$ and $\tilde{T}_{R}$ ($\tilde{d}_{R}$ and $\tilde{B}_{R}$) are a mixture of $u_{R}$ and $T_{R}$ ($d_{R}$ and $B_{R}$) with opposite $U(1)$ charges. The same consideration applies for left-handed components of these quarks, $\tilde{u}_{L}$ and $\tilde{T}_{L}$ ($\tilde{d}_{L}$ and $\tilde{B}_{L}$). In fact, in the case of left-handed components, we should also consider that while $u_{L}$ and $d_{L}$ form a $SU(2)$ doublet, $T_{L}$ and $B_{L}$ do not. Putting these considerations together, we find out that the flavor changing neutral currents with vertices including third and/or fourth generations are forbidden, implying that Kaons, $D$-mesons and $B$-mesons will behave as the SM predicts but we can have vertices such as
$\bar{\tilde{B}}_{L}\gamma^{\mu}\tilde{d}_{L}Z^{\prime}_{\mu}\ ,\ \bar{\tilde{B}}_{L}\gamma^{\mu}\tilde{d}_{L}Z_{\mu}\ ,\ \bar{\tilde{T}}_{L}\gamma^{\mu}\tilde{u}_{L}Z^{\prime}_{\mu}\ {\rm and}\ \bar{\tilde{T}}_{L}\gamma^{\mu}\tilde{u}_{L}Z_{\mu}$ (20)
given by the relevant gauge coupling times the corresponding mixing angle. Similarly, we have vertices
$\bar{\tilde{B}}_{R}\gamma^{\mu}\tilde{d}_{R}Z^{\prime}_{\mu}\ {\rm and}\ \bar{\tilde{T}}_{R}\gamma^{\mu}\tilde{u}_{R}Z^{\prime}_{\mu}\ .$
Remembering that $u_{L}$ and $d_{L}$ are defined to be perpendicular to mass eigenstates $\tilde{c}_{R}$, $\tilde{t}_{R}$, $\tilde{s}_{R}$ and $\tilde{b}_{R}$, we can therefore write the charged current interactions as
$W_{\mu}(\bar{u}_{L}\ \bar{\tilde{c}}_{L}\ \bar{\tilde{t}}_{L})\gamma^{\mu}V_{CKM}\left(\begin{matrix}d_{L}\cr\tilde{s}_{L}\cr\tilde{b}_{L}\end{matrix}\right)+H.c.=$ (21)
$W_{\mu}(\bar{\tilde{u}}_{L}\cos\alpha+\bar{\tilde{T}}_{L}\sin\alpha\ \ \bar{\tilde{c}}_{L}\ \bar{\tilde{t}}_{L})\gamma^{\mu}V_{CKM}\left(\begin{matrix}\tilde{d}_{L}\cos\beta+\tilde{B}_{L}\sin\beta\cr\tilde{s}_{L}\cr\tilde{b}_{L}\end{matrix}\right)+H.c.$
where $\sin\alpha\simeq\lambda_{Tu}\langle H\rangle/M_{T}$ and $\sin\beta\simeq\lambda_{Bd}\langle H\rangle/M_{B}$. As mentioned in sect. II, the dominant decay modes of $\tilde{B}$ and $\tilde{T}$ are $\tilde{B}\to W\tilde{u}$, $\tilde{B}\to Z\tilde{d}$, $\tilde{T}\to W\tilde{d}$ and $\tilde{T}\to Z\tilde{u}$. Remember that $\epsilon^{Au}_{\tau\tau}=\epsilon^{Ad}_{\tau\tau}$ is suppressed by the inverse of $\langle\phi\rangle^{2}=M_{T}^{2}/\lambda_{T}^{2}=M_{B}^{2}/\lambda_{B}^{2}$. Thus, to have sizable NSI, the masses of $T$ and $B$ should be close to the present lower bound. $B$ and $T$ can therefore be pair produced via QCD interaction at the LHC and detected as jet+$W$ or jet+$Z$. As a result, there is a good chance of discovery of these particles in near future.
Notice that the unitarity of the $3\times 3$ light quark mixing will be violated by $\cos\alpha\cos\beta$. Taking $\lambda_{Tu}\langle H\rangle\sim M_{Tu}\sim\sqrt{m_{u}M_{T}}$ and $\lambda_{Bd}\langle H\rangle\sim M_{Bd}\sim\sqrt{m_{d}M_{B}}$, we find $1-\cos\alpha\cos\beta\sim 3\times 10^{-6}$ which is too small to be tested by the unitarity of the CKM matrix.
IV Possible generalization of the toy model
In sect. II, we presented a model for $\epsilon_{\tau\tau}^{Au}=\epsilon^{Ad}_{\tau\tau}$. The reason why we have focused on the $\tau\tau$ component is that this component is less experimentally restricted. We could also assign a $U(1)$ charge to the second generation of leptons and obtain $\epsilon_{\mu\mu}^{Au}=\epsilon^{Ad}_{\mu\mu}\neq 0$. As long as the $U(1)$ charges of the second and third generation leptons summed up to 9, the $U(1)-SU(2)-SU(2)$ anomaly would be canceled. Exploring the parameter space is beyond the scope of this paper. Similarly to the models for vectorial NSI presented in Farzan (2020); Farzan and Heeck (2016), arbitrary flavor structure can be obtained by adding sterile neutrinos with a $U(1)$ charge mixed with the active neutrinos.
To cancel the $U(1)-SU(3)-SU(3)$ anomaly, we have added a pair of quarks. The lower bound on the masses of these colored particles leads to an upper bound on $\epsilon^{Au}=\epsilon^{Ad}$. Instead of introducing new quarks, we could assign $U(1)$ charges to $s_{R}$ and $c_{R}$, $\alpha_{s}$ and $\alpha_{c}$, such that $\alpha_{s}+\alpha_{c}=6$. This requires new Higgs doublets with suitable $U(1)$ charges to give mass to the first and second quarks. Then, to cancel various $U(1)$ and $U_{Y}(1)$ anomalies, new charged particles (charged both under $U_{em}(1)$ and new $U(1)$) should be added. In the minimal version, with addition of only a pair of chiral fermions, $E_{L}$ and $E_{R}$, with electric charge of 1 and different $U(1)$ charges, the anomaly cancellation requires fractional $U(1)$ charges. The mass of $E$ should be obtained after breaking of the $U(1)$ symmetry via a Yukawa coupling of form $\phi E_{L}^{\dagger}E_{R}$. The new fermion is colorless so the the lower bounds on its mass is less stringent. As a result, $\langle\phi\rangle$ can be smaller than the case with colorful new fermions. Thus, $\epsilon^{Au}=\epsilon^{Ad}$ can be larger in this version of the model without new particles.
In sect. II, we imposed a $Z_{2}$ symmetry to make $\chi_{L}$ stable and a suitable dark matter candidate. If future direct dark matter searches do not find it (or more precisely, for a given value of $\epsilon^{Au}_{\tau\tau}\sim 1$, the direct dark matter searches restrict the $\chi$ mass to below 100 MeV), we may relax the $Z_{2}$ symmetry and allow its decay. The dark matter should then find another candidate.
V Summary and discussion
[…]
To obtain axial coupling, the $U(1)$ charges of right-handed and left-handed first generation quarks have to be opposite which leads to a $U(1)-SU(3)-SU(3)$ anomaly. To cancel the anomaly, we introduce new quarks with masses of $\sim 1500$ GeV. In our model, the mass generation mechanism for the light quarks is not Yukawa coupling to the SM Higgs but mixings with new heavy quarks. In other words, there will be a seesaw mechanism, explaining the smallness of the $u$ and $d$ quark masses relative to the top mass. The heavy quarks can be pair produced at the hadron colliders and can then decay to the first generation quarks plus $W$ or $Z$ gauge boson. Being only slightly heavier than the present lower bound, there is a bright prospect of discovery of the heavy quarks at the near future by the LHC.
Finally, to cancel the $U(1)$ anomaly, a SM singlet chiral fermion, $\chi_{L}$, has to be added. Imposing a $Z_{2}$ symmetry, $\chi_{L}$ can be made stable and therefore a suitable dark matter particle. The dark matter can undergo spin dependent scattering off quarks. The bounds from spin dependent direct dark matter searches already push the $\chi_{L}$ mass to be smaller than 2 GeV. The future spin dependent dark matter experiments can discover $\chi_{L}$. We have discussed the $\chi_{L}$ production mechanism in the early universe and pointed out three alternative scenarios to obtain the observed dark matter abundance. In our model, the indirect dark matter searches are expected to yield null result and the dark matter will act as collision-less at structure formation as well as at collision of the galaxy clusters.
In summary, our model is super-testable with four potential discoveries around the corner: (1) The impact of $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$ on the number of neutral current events at the far detector of DUNE; (2) discovery of $Z^{\prime}$ at the high luminosity LHC; (3) discovery of heavy quarks at the high luminosity LHC and (4) discovery of $\sim$ GeV mass dark matter at spin dependent direct dark matter search experiments.
We have discussed a few possible generalizations of the model. In particular, we have discussed how an arbitrary lepton flavor structure for $\epsilon^{Au}=\epsilon^{Ad}$ can be obtained. We have also discussed a variation of the model without heavy quarks but with new charged singlets.
Acknowledgements
This work has partially been supported by the European Unions Framework Programme for Research and Innovation Horizon 2020 under grant H2020-MSCA-ITN2019/860881-HIDDeN as well as under the Marie Sklodowska-Curie Staff Exchange grant agreement No 5 101086085-ASYMMETRY. YF would like to acknowledge support from ICTP through the Associates Programme and from the Simons Foundation through grant number 284558FY19. YF thanks ICTP high energy group for their hospitality during her visit when this paper was written.